It was found out that average intake of medicine in a group of patients is 110mg with standard deviation 10mg per week.
(a) a patient is chosen at random. Find the probability that his/her intake is in betweem 100mg and 115mg per week.
(b) a sample of 3 patients is chosen at random. Find the probability that the sample mean (average) intake exceeds 115mg per week.

Answer by stanbon(75887)

It was found out that average intake of medicine in a group of patients is 110mg with standard deviation 10mg per week.
-------------------------------
(a) a patient is chosen at random. Find the probability that his/her intake is in betweem 100mg and 115mg per week.
z(100) = (100-110)/10 = -1
z(115) = (115-110)/10 = 1/2
P(100 < x < 115) = P(-1< z <1/2) = 0.5328
------------------------------------------
(b) a sample of 3 patients is chosen at random. Find the probability that the sample mean (average) intake exceeds 115mg per week.
t(115) = (115-110)/(10/sqrt(3)) = 5/(10/sqrt(3)) = 0.8660
---
P(x > 115) = P(t > 0.8660 when df = 2) = 0.2388
==================
Cheers,
Stan H.
==============